A taxable bond with a coupon rate of 5.00% has a market price of 98.29% of par. The bond matures in 16.00 years ans pays semi-annually. Assume an investor has a 16.00% marginal tax rate. The investor would prefer otherwise identical tax-exempt bond if it's yield to maturity was more than _____% (Round to 2 decimal places)
Answer #1

Let Par Value =1000
Price of Bond =98.29%*1000 =982.90
Semi annual Coupon =5%*1000/2 =25
Number of Periods =16*2 =32
YTM or cost of Debt using financial calculator:
N=32;PMT=25;PV=-982.90;FV=1000;CPT I/Y =2.583%
YTM =2*2.583% =5.1667%

The YTM for tax exempt bond =5.1667%/(1-tax Rate) =5.1667%/(1-16%) =6.15%